For this function, calculate the equation of the tangent line at the given point:
f(x) = sqrt of 6x at the point (6,6).

Answer by tommyt3rd(5050)

details posted elsewhere...
y'(6)=3/sqrt(36)=1/2
y-6=1/2(x-6)